# Simulating Human-like Daily Activities with Desire-driven Autonomy

### v2

@@ 1 Introduction
Figure 1: Motivation Behind Different Activity Generation Approaches. Our Desire-driven Autonomous Agent (D2A) follows the ”Act to Satisfy Intrinsic Desires” principle (right), whereas most existing methods or frameworks focus on goal reasoning or characteristic-driven actions.
@@ 5.1 Desire Generation
The theory of needs suggests that individual behavior is motivated by meeting the individual’s needs or wants, such as safety, social, esteem, and self-actualization. Furthermore, individuals may prioritize different needs based on their unique profiles, traits, and personalities. These considerations underscore the complexity of human needs and illustrate how these needs drive behavior. To simulate human behavior, we developed a value system that maps personality profile $p$ with the expected desire values $v^{*}$ (See Appendix E and F). Specifically, for the indoor, single-agent environment, we define 11 dimensions of desired value. The physiological dimension, representing fundamental survival needs, is further divided into categories such as ‘hunger’, ‘thirst’, ‘sleepiness’, ‘cleanliness’, ‘comfort’, and ‘health’. Safety needs are encapsulated by the ‘safeness’ dimension. To capture love and belonging needs, which reflect the desire for relationships, we introduce the ‘social connectivity’ dimension. Finally, the self-actualization needs, linked to personal growth, are categorized into ‘joy’, ‘passion’, and ‘spiritual satisfaction’. For the outdoor party multi-agents environment, we further introduce ‘recognition’, ‘sense of control’, and ‘sense of superiority’ three desire dimensions to steer activity generation with higher-level motivations in a social environment. To quantitatively track changes in the level of these desires and qualitatively translate numerical values into descriptive states, we apply a [0-10] Likert scale for each dimension, mapping numerical values into textual descriptive outcomes (See Appendix G).
@@
[…]
For instance, the expected social connectivity value for extremely sociable Alice is 9, while for slightly sociable Alice, it is 7. During the agent’s initialization, each adjective is randomly paired with an adverb of degree (See Appendix E). After this automatic mapping, each desired dimension is assigned its expected value $v_{d}^{*}$. The initial value $v_{d}^{0}$ for each desire dimension is randomly selected from the range $[0,10]$, completing the desire generation process for our D2A framework.
@@ 5.2 Value System for Desire Evaluation
As shown in Figure 2, the Value System executes three key procedures at each simulation step: Quantitative Value Deduction, Qualitative Value Description, and Value Update. The Quantitative Value Deduction procedure models the fluctuating nature of human desires by incorporating a decay mechanism for all intrinsic values. At the beginning of each time step $t$, the agent’s intrinsic value $v_{d}^{t-1}$ for desire dimension $d$ is reduced by 1 with a probability $p_{d}$. This probability is determined by the agent’s profile characteristics, following the same mechanism used for the expected desire value mapping described in Appendix E and F. Since fulfilling these desires serves as D2A’s intrinsic motivation, the Quantitative Value Deduction procedure drives the agent to continuously generate desire-driven activities.
Since many studies have shown that large language models often struggle with interpreting numerical values, we designed the Qualitative Value Description procedure to translate these values into meaningful textual desire descriptions. As shown in our ablation studies (Sec. 6.4), the descriptions are important for the agent to understand its current desire states so as to generate appropriate activities. To achieve this, we implemented a workflow that converts all intrinsic numerical values $v^{t}$ into descriptive desire sentences $D^{t}$ by asking a set of predefined questions (Appendix H) to backbone LLMs at each simulation step $t$. The agent is expected to recognize the shortfall $\delta^{t}$ between its current desires and the expected values, based on the descriptive desire $D^{t}$ and its internal profile $p$ throughout the simulation.
The Desire-driven Planner then takes over and generates the activity $a_{t}$. The environment subsequently provides the agent with the corresponding observation response $o_{t}$. Afterward, the Value System initiates the Value Update procedure, updating the intrinsic numerical values $v^{t}$ of the desire components based on the activity $a_{t}$, the observation $o_{t}$, the previous values $v^{t-1}$, and the value description prompts (Appendix G).
@@ 5.3 Desire-Driven Planner
Inspired by the work of Tree of Thoughts (Yao et al., 2024), we argue that the paradigm of presenting multiple candidate activities and evaluating them enables the agent to select the best action. Thus, during the Activity Proposal procedure (Appendix J.2), we prompt the agent to generate $N$ candidates (also referred to as planner width, which is set to 3 in our default experimental configuration) activities that may have positive impacts on its desires given $p$, $e$, $a_{0:t-1}$, $o_{0:t-1}$, and desire descriptions $D^{t}$:
@@
These $N$ candidate activities are then passed on to the Activity Evaluation procedure (Appendix J.3). The Planner evaluates each activity $a_{t}^{i}$ by imagining the resulting desire states $\hat{D_{t}^{i}}$ that the agent would experience after taking action $a_{t}^{i}$:
@@
Finally, in the Activity Selection procedure (Appendix J.4), the agent compares the forecasted desire states $\hat{D_{t}^{i}}$ for all $N$ proposed actions. Based on these evaluations, the Planner selects the activity $a_{t}$ that is expected to best fulfill the agent’s intrinsic desires:
@@ 6.1 Environment
We compare our Desire-Driven Agent with 3 baselines: ReAct (Yao et al. (2023)) agent reasons before taking actions; LLMob (Wang et al. (2024a)) generates activity plans by extracting motivations from the profile and producing the most likely activities accordingly; BabyAGI (Nakajima (2023)) maintains a prioritized task list and selects the next activity based on priority. To evaluate the effectiveness of the specific designs made for D2A, desire states $D^{t}$ are not provided to the baseline agents. Instead, they are given the goal, profile, or more descriptive habits (Appendix K) generated by the profile according to their respective methods. As a reference for human behavior, we also include a human-controlled agent, where the player is given the same desired value provided to the D2A. For all four agent-based methods, we used LLaMA3.1-70B as the default backbone model for both the agents and the environment controller.
@@ 6.2 Evaluation Metrics
measures how closely an agent’s behavior resembles that of a human. Given the difficulty of directly measuring human likeness, we define likelihood by employing an external model to assess the win rate between activity sequences generated by two agents, determining which sequence appears more human-like. We conduct this ”Player vs Player” comparison across the activity sequences produced by the baselines and the D2A agent. Specifically, after obtaining all action sequences $A_{p}$ for agent $p$, represented as $[A^{1}_{p},\ldots,A^{N}_{p}]$, we randomly select $seq_{i}$, $seq_{j}$ from $[A_{i}]$, $[A_{j}]$, respectively, for a given pair of agents $i$, $j$. These sequences are then compared using GPT-4o. This process is repeated 100 times for each agent pair $i$ and $j$ ($\forall i\neq j$), and the win rate for each agent is calculated. The results are visualized using a heatmap.
@@ 6.3.1 Does D2A Generate More Human-like Daily Activities?
We conducted 15 trials for each agent under identical initialization settings across all experiments. After gathering each agent’s activity sequences, we found differences in output styles (See Appendix U) that could influence the evaluation by GPT-4o. To mitigate this, we preprocessed all sequences using Llama3.1 70B for style-level rewriting. This ensured more consistent activity sequences across agents, allowing for a fair comparison in the subsequent evaluation.
We used the method mentioned in Section (6.2) to obtain the relative likelihood map of approaches. When evaluating the likelihood of the activity sequences, GPT-4o is instructed to take the following criteria into account: Naturalness: How well the activity aligns with the person’s innate abilities, habits, and environment. Coherence: How logically and seamlessly different actions or steps within a sequence fit together to achieve a goal. Plausibility: How reasonable, likely, or believable a sequence of actions is, given the circumstances, context, and known behavior patterns. GPT-4o was asked to consider all three criteria together and output a final determination of which sequence was more human-like. See Appendix W.1 for the prompt.
@@
The results presented in Figure 3 indicate that the D2A agent outperforms other baseline agents. We validated through human experiments that GPT-4o’s evaluations exhibit a high level of consistency with human judges (See Appendix 1). This superior performance may be attributed to the fact that D2A is motivated by desire, aligning with the theory of human needs. Guided by desire, D2A effectively simulates human-like activity sequences, demonstrating a level of naturalness because of the recognizable characteristics of its profile. Furthermore, coherence is maintained as D2A considers its current desires based on the activity proposal component, continuously engaging in specific actions to achieve its goal of living comfortably until its desires are fulfilled. In contrast, ReAct, as the foundational agent, operates on a goal-reasoning framework. Its reliance on reasoning for action selection ensures reasonable interactions with environmental elements, thereby enhancing plausibility. However, the absence of a structured plan diminishes both the coherence and naturalness of its actions. BabyAGI, on the other hand, generates subsequent tasks based on its profile and motivational factors and maintains the priority list based on the observation. Although this approach facilitates coherence, it limits the diversity, as each activity is based on the outcomes of the previous one. Consequently, the agent’s tendency to engage repetitively in similar behaviors (enhancing its hunger based on the profile (See Appendix R), may result in diminished performance compared to D2A due to a lack of action diversity. LLMob leverages both planning and motivation derived from its profile to create a broad activity space. However, since it is characteristic-oriented, it keeps performing activities related to the habits derived from its profile. As a result, there may be no connection between activities, reducing overall coherence. The comparison of action diversity can be found in Appendix D.1.
@@ 6.3.3 Can D2A Generalize to Social Interactions Intensive Environment?
To assess the adaptability of our D2A framework within complex social interaction scenarios, we conducted experiments in the outdoor environment. The results, presented in Figure 5, reveal that D2A effectively minimizes dissatisfaction across desire dimensions, mirroring its performance in the indoor settings. This consistent behavior underscores the framework’s robust generalization capabilities across different environments. Detailed simulated activity sequences from the outdoor environment are available in Appendix V. Notably, in this setting, D2A autonomously interacts with non-player characters (NPCs) through conversations and games, ultimately reducing intrinsic dissatisfaction. Besides, as illustrated in Figure 13, D2A demonstrates the ability to generate activities that are well-balanced across diverse categories.
@@ 6.4 Ablation Study
As described in Section 5.3, our D2A does not contain a multi-step planning module that plans the next few steps’ activities ahead, like some other approaches in consideration of the dynamic nature of desires. We performed an ablation study by adding Concordia’s multi-step planning component, which updates the plan for the next few hours based on current desired states. As shown in Figure 6(c), the absence of the multi-step planning component allows D2A to exhibit a better ability to respond to immediate desire states. This may be because when the multi-step planning component is present, D2A tends to adhere to the pre-defined activities in the plan, which may limit its responsiveness to real-time changes in the desired value. Furthermore, the plan is scheduled in hourly increments, and D2A cannot effectively predict which desires will experience the greatest shortfall in the next several hours. This limitation may contribute to the slight decline in performance observed when the plan component is included. We also validated the adaptability of our framework across different backbone large language models, the details can be found in Appendix O.
@@ A.3 Background Agents in Outdoor Environment
Goal: ”Bob hopes to meet the guests’ needs.”,
@@
03 Jul 2007 00:00:00 When Bob was 23 years old, he was hired as a staff member at Central Park’s ”Enjoy Your Life” party event, and he quickly became a favorite among attendees and colleagues alike, known for his infectious energy and ability to connect with anyone. As he stood in the networking lounge, surrounded by people laughing and chatting, Bob felt a sense of belonging and purpose that he had never felt before. This experience was the culmination of all his hard work and dedication, and Bob knew that he had finally found his dream job.
@@
03 Jul 2005 00:00:00 At 21 years old, Charlie stumbled upon Central Park’s ”Enjoy Your Life” party event, and he was immediately drawn to the vibrant atmosphere and eclectic crowd, feeling like he had found his tribe, his people. He spent hours taking photos and networking with other attendees, and for the first time, he felt like he was part of a community that shared his passions and values. This experience marked the beginning of a new chapter in Charlie’s life, as he began to prioritize his own happiness and well-being.
@@ Appendix B Consistency between Human Annotators and GPT-4o Evaluations
To validate the robustness of GPT-4o’s evaluations, we randomly selected 19 pairs of activity sequences for pairwise comparison (16 pairs from the indoor scenarios and 3 pairs from the outdoor scenarios) and created three questionnaires, each assigned to 8 human annotators. The human annotators were asked to judge which sequence in each pair appeared more human-like. Based on the level of agreement among the annotators, we categorized the 18 samples into three groups: over 75% agreement on one side—high consensus among humans; 50.1%–74.9%—moderate human preference; and exactly 50%—indicating humans found it difficult to distinguish between the two. We also provided the same pairs to GPT-4o, using our evaluation promptW.1 and both sequences as context, and asked it to determine which sequence appeared more human-like. The consistency rate between human evaluations and GPT-4o’s assessments is summarized as follows.
@@
High consensus ($>75\%$ agreement) 11/19 100

### v3

@@ 1 Introduction
Figure 1: Motivation Behind Different Activity Generation Approaches. Our Desire-driven Autonomous Agent (D2A) follows the “Act to Satisfy Intrinsic Desires” principle (right), whereas most existing methods or frameworks focus on goal reasoning or characteristic-driven actions.
@@ 5.1 Desire Generation
The theory of needs suggests that individual behavior is motivated by meeting the individual’s needs or wants, such as safety, social, esteem, and self-actualization. Furthermore, individuals may prioritize different needs based on their unique profiles, traits, and personalities. These considerations underscore the complexity of human needs and illustrate how these needs drive behavior. To simulate human behavior, we developed a value system that maps personality profile $p$ with the expected desire values $v^{*}$ (See Appendix F and G). Specifically, for the indoor, single-agent environment, we define 11 dimensions of desired value. The physiological dimension, representing fundamental survival needs, is further divided into categories such as ‘hunger’, ‘thirst’, ‘sleepiness’, ‘cleanliness’, ‘comfort’, and ‘health’. Safety needs are encapsulated by the ‘safeness’ dimension. To capture love and belonging needs, which reflect the desire for relationships, we introduce the ‘social connectivity’ dimension. Finally, the self-actualization needs, linked to personal growth, are categorized into ‘joy’, ‘passion’, and ‘spiritual satisfaction’. For the outdoor party multi-agents environment, we further introduce ‘recognition’, ‘sense of control’, and ‘sense of superiority’ three desire dimensions to steer activity generation with higher-level motivations in a social environment. To quantitatively track changes in the level of these desires and qualitatively translate numerical values into descriptive states, we apply a [0-10] Likert scale for each dimension, mapping numerical values into textual descriptive outcomes (See Appendix H).
@@
$\{\text{degree, adjective}\}\xrightarrow{\text{predefined map}}v_{d}^{*}$ (2)
For instance, the expected social connectivity value for extremely sociable Alice is 9, while for slightly sociable Alice, it is 7. During the agent’s initialization, each adjective is randomly paired with an adverb of degree (See Appendix F). After this automatic mapping, each desired dimension is assigned its expected value $v_{d}^{*}$. The initial value $v_{d}^{0}$ for each desire dimension is randomly selected from the range $[0,10]$, completing the desire generation process for our D2A framework.
@@ 5.2 Value System for Desire Evaluation
As shown in Figure 2, the Value System executes three key procedures at each simulation step: Quantitative Value Deduction, Qualitative Value Description, and Value Update. The Quantitative Value Deduction procedure models the fluctuating nature of human desires by incorporating a decay mechanism for all intrinsic values. At the beginning of each time step $t$, the agent’s intrinsic value $v_{d}^{t-1}$ for desire dimension $d$ is reduced by 1 with a probability $p_{d}$. This probability is determined by the agent’s profile characteristics, following the same mechanism used for the expected desire value mapping described in Appendix F and G. Since fulfilling these desires serves as D2A’s intrinsic motivation, the Quantitative Value Deduction procedure drives the agent to continuously generate desire-driven activities.
Since many studies have shown that large language models often struggle with interpreting numerical values, we designed the Qualitative Value Description procedure to translate these values into meaningful textual desire descriptions. As shown in our ablation studies (Sec. 6.4), the descriptions are important for the agent to understand its current desire states so as to generate appropriate activities. To achieve this, we implemented a workflow that converts all intrinsic numerical values $v^{t}$ into descriptive desire sentences $D^{t}$ by asking a set of predefined questions (Appendix I) to backbone LLMs at each simulation step $t$. The agent is expected to recognize the shortfall $\delta^{t}$ between its current desires and the expected values, based on the descriptive desire $D^{t}$ and its internal profile $p$ throughout the simulation.
The Desire-driven Planner then takes over and generates the activity $a_{t}$. The environment subsequently provides the agent with the corresponding observation response $o_{t}$. Afterward, the Value System initiates the Value Update procedure, updating the intrinsic numerical values $v^{t}$ of the desire components based on the activity $a_{t}$, the observation $o_{t}$, the previous values $v^{t-1}$, and the value description prompts (Appendix H).
@@ 5.3 Desire-Driven Planner
Inspired by the work of Tree of Thoughts (Yao et al., 2024), we argue that the paradigm of presenting multiple candidate activities and evaluating them enables the agent to select the best action. Thus, during the Activity Proposal procedure (Appendix K.2), we prompt the agent to generate $N$ candidates (also referred to as planner width, which is set to 3 in our default experimental configuration) activities that may have positive impacts on its desires given $p$, $e$, $a_{0:t-1}$, $o_{0:t-1}$, and desire descriptions $D^{t}$:
@@
These $N$ candidate activities are then passed on to the Activity Evaluation procedure (Appendix K.3). The Planner evaluates each activity $a_{t}^{i}$ by imagining the resulting desire states $\hat{D_{t}^{i}}$ that the agent would experience after taking action $a_{t}^{i}$:
@@
Finally, in the Activity Selection procedure (Appendix K.4), the agent compares the forecasted desire states $\hat{D_{t}^{i}}$ for all $N$ proposed actions. Based on these evaluations, the Planner selects the activity $a_{t}$ that is expected to best fulfill the agent’s intrinsic desires:
@@ 6.1 Environment
We compare our Desire-Driven Agent with 3 baselines: ReAct (Yao et al. (2023)) agent reasons before taking actions; LLMob (Wang et al. (2024a)) generates activity plans by extracting motivations from the profile and producing the most likely activities accordingly; BabyAGI (Nakajima (2023)) maintains a prioritized task list and selects the next activity based on priority. To evaluate the effectiveness of the specific designs made for D2A, desire states $D^{t}$ are not provided to the baseline agents. Instead, they are given the goal, profile, or more descriptive habits (Appendix L) generated by the profile according to their respective methods. As a reference for human behavior, we also include a human-controlled agent, where the player is given the same desired value provided to the D2A. For all four agent-based methods, we used LLaMA3.1-70B as the default backbone model for both the agents and the environment controller.
@@ 6.2 Evaluation Metrics
measures how closely an agent’s behavior resembles that of a human. Given the difficulty of directly measuring human likeness, we define likelihood by employing an external model to assess the win rate between activity sequences generated by two agents, determining which sequence appears more human-like. We conduct this “Player vs Player” comparison across the activity sequences produced by the baselines and the D2A agent. Specifically, after obtaining all action sequences $A_{p}$ for agent $p$, represented as $A^{1}_{p},\ldots,A^{N}_{p}]$, we randomly select $seq_{i}$, $seq_{j}$ from $A_{i}]$, $A_{j}]$, respectively, for a given pair of agents $i$, $j$. These sequences are then compared using GPT-4o. This process is repeated 100 times for each agent pair $i$ and $j$ ($\forall i\neq j$), and the win rate for each agent is calculated. The results are visualized using a heatmap.
@@ 6.3.1 Does D2A Generate More Human-like Daily Activities?
We conducted 15 trials for each agent under identical initialization settings across all experiments. After gathering each agent’s activity sequences, we found differences in output styles (See Appendix V) that could influence the evaluation by GPT-4o. To mitigate this, we preprocessed all sequences using Llama3.1 70B for style-level rewriting. This ensured more consistent activity sequences across agents, allowing for a fair comparison in the subsequent evaluation.
We used the method mentioned in Section (6.2) to obtain the relative likelihood map of approaches. When evaluating the likelihood of the activity sequences, GPT-4o is instructed to take the following criteria into account: Naturalness: How well the activity aligns with the person’s innate abilities, habits, and environment. Coherence: How logically and seamlessly different actions or steps within a sequence fit together to achieve a goal. Plausibility: How reasonable, likely, or believable a sequence of actions is, given the circumstances, context, and known behavior patterns. GPT-4o was asked to consider all three criteria together and output a final determination of which sequence was more human-like. See Appendix X.1 for the prompt.
@@
The results in Figure 3 indicate that the D2A agent outperforms other baseline agents. We validated through human experiments that GPT-4o’s evaluations exhibit a high level of consistency with human judges (See Appendix 1). This superior performance may be attributed to the fact that D2A is motivated by desire, aligning with the theory of human needs. Guided by desire, D2A effectively simulates human-like activity sequences, demonstrating a level of naturalness because of the recognizable characteristics of its profile. Furthermore, coherence is maintained as D2A considers its current desires based on the activity proposal component, continuously engaging in specific actions to achieve its goal of living comfortably until its desires are fulfilled. In contrast, ReAct, as the foundational agent, operates on a goal-reasoning framework. Its reliance on reasoning for action selection ensures reasonable interactions with environmental elements, thereby enhancing plausibility. However, the absence of a structured plan diminishes both the coherence and naturalness of its actions. BabyAGI, on the other hand, generates subsequent tasks based on its profile and motivational factors and maintains the priority list based on the observation. Although this approach facilitates coherence, it limits the diversity, as each activity is based on the outcomes of the previous one. Consequently, the agent’s tendency to engage repetitively in similar behaviors (enhancing its hunger based on the profile (See Appendix S), may result in diminished performance compared to D2A due to a lack of action diversity. LLMob leverages both planning and motivation derived from its profile to create a broad activity space. However, since it is characteristic-oriented, it keeps performing activities related to the habits derived from its profile. As a result, there may be no connection between activities, reducing overall coherence. The comparison of action diversity can be found in Appendix D.1.
@@ 6.3.3 Can D2A Generalize to Social Interactions Intensive Environment?
To assess the adaptability of our D2A framework within complex social interaction scenarios, we conducted experiments in the outdoor environment. The results, presented in Figure 5, reveal that D2A effectively minimizes dissatisfaction across desire dimensions, mirroring its performance in the indoor settings. This consistent behavior underscores the framework’s robust generalization capabilities across different environments. Detailed simulated activity sequences from the outdoor environment are available in Appendix W. Notably, in this setting, D2A autonomously interacts with non-player characters (NPCs) through conversations and games, ultimately reducing intrinsic dissatisfaction. Besides, as illustrated in Figure 12, D2A demonstrates the ability to generate activities that are well-balanced across diverse categories.
@@ 6.4 Ablation Study
As described in Section 5.3, our D2A does not contain a multi-step planning module that plans the next few steps’ activities ahead, like some other approaches in consideration of the dynamic nature of desires. We performed an ablation study by adding Concordia’s multi-step planning component, which updates the plan for the next few hours based on current desired states. As shown in Figure 6(c), the absence of the multi-step planning component allows D2A to exhibit a better ability to respond to immediate desire states. This may be because when the multi-step planning component is present, D2A tends to adhere to the pre-defined activities in the plan, which may limit its responsiveness to real-time changes in the desired value. Furthermore, the plan is scheduled in hourly increments, and D2A cannot effectively predict which desires will experience the greatest shortfall in the next several hours. This limitation may contribute to the slight decline in performance observed when the plan component is included. We also validated the adaptability of our framework across different backbone large language models, the details can be found in Appendix P.
@@ A.3 Background Agents in Outdoor Environment
Goal: “Bob hopes to meet the guests’ needs”,
@@
03 Jul 2007 00:00:00 When Bob was 23 years old, he was hired as a staff member at Central Park’s “Enjoy Your Life” party event, and he quickly became a favorite among attendees and colleagues alike, known for his infectious energy and ability to connect with anyone. As he stood in the networking lounge, surrounded by people laughing and chatting, Bob felt a sense of belonging and purpose that he had never felt before. This experience was the culmination of all his hard work and dedication, and Bob knew that he had finally found his dream job.
@@
03 Jul 2005 00:00:00 At 21 years old, Charlie stumbled upon Central Park’s “Enjoy Your Life” party event, and he was immediately drawn to the vibrant atmosphere and eclectic crowd, feeling like he had found his tribe, his people. He spent hours taking photos and networking with other attendees, and for the first time, he felt like he was part of a community that shared his passions and values. This experience marked the beginning of a new chapter in Charlie’s life, as he began to prioritize his own happiness and well-being.
@@ Appendix B Consistency between Human Annotators and GPT-4o Evaluations
To validate the robustness of GPT-4o’s evaluations, we randomly selected 19 pairs of activity sequences for pairwise comparison (16 pairs from the indoor scenarios and 3 pairs from the outdoor scenarios) and created three questionnaires, each assigned to 8 human annotators. The human annotators were asked to judge which sequence in each pair appeared more human-like. Based on the level of agreement among the annotators, we categorized the 18 samples into three groups: over 75% agreement on one side—high consensus among humans; 50.1%–74.9%—moderate human preference; and exactly 50%—indicating humans found it difficult to distinguish between the two. We also provided the same pairs to GPT-4o, using our evaluation promptX.1 and both sequences as context, and asked it to determine which sequence appeared more human-like. The consistency rate between human evaluations and GPT-4o’s assessments is summarized as follows.
@@
High consensus ($75\%$ agreement) 11/19 100